### v2

# Homological Filling and Minimal Varifolds in Four-Dimensional Einstein Manifolds

Abstract.
We study the smallest area $A_{\min}(M,g)$ of a $2$-dimensional stationary integral varifold in a closed Einstein $4$-manifold $(M^{4},g)$ with
$\operatorname{Ric}_{g}=\lambda g,\quad|\lambda|\leq 3,\quad\operatorname{Vol}(M,g)\geq v>0,\quad\operatorname{diam}(M,g)\leq D,\quad H_{1}(M;\mathbb{Z})=0.$
Building on the previous work [24] on homological filling functions, we show that for every $(M^{4},g)$ in this Einstein class there is an upper bound
$A_{\min}(M,g)\leq F_{\operatorname{Ein}}(v,D),$
where $F_{\operatorname{Ein}}$ depends only on $(v,D)$. The proof uses the Einstein-specific analytic input of a uniform Sobolev inequality, a global $L^{2}$-curvature bound, and standard $\varepsilon$-regularity/harmonic-radius estimates.
1. Introduction
Let $(M^{4},g)$ be a closed Riemannian manifold and denote by $A_{\min}(M,g)$ the area of a smallest $2$-dimensional stationary integral varifold in $(M,g)$. In general, the existence of such a minimal object follows from the Almgren–Pitts min–max theory [21], and in low codimension and low dimension one can often upgrade the varifold to a smooth embedded minimal hypersurface using the regularity theory of Schoen–Simon–Yau [23, 22].
In this article we are interested in bounding $A_{\min}(M,g)$ in terms of global geometric data in the special case when $(M,g)$ is an Einstein $4$-manifold subject to natural non-collapsing and diameter bounds. Throughout we assume that
(1) $\operatorname{Ric}_{g}=\lambda g,\quad|\lambda|\leq 3,\quad\operatorname{Vol}(M,g)\geq v>0,\quad\operatorname{diam}(M,g)\leq D,\quad H_{1}(M;\mathbb{Z})=0.$
We write $\mathcal{E}_{1}(4,v,D)$ for the class of all closed Einstein $4$-manifolds satisfying (1).
This paper is a sequel to our previous work [24], where we considered the larger class with $|{\rm Ric}_{g}|\leq 3,\ \operatorname{Vol}(M,g)\geq v>0,\operatorname{diam}(M,g)\leq D,\ H_{1}(M;\mathbb{Z})=0,$ and proved an affine upper bound for the first homological filling function $\operatorname{FH}_{1}$ of $(M,g)$. More precisely, for every $(M^{4},g)\in\mathcal{M}_{1}(4,v,D)$ we showed that there exist functions $f_{1}(v,D),f_{2}(v,D)>0$ such that
$\operatorname{FH}_{1}(l)\ \leq\ f_{1}(v,D)\,l+f_{2}(v,D)\qquad\text{for all }l\geq 0.$
Combined with an effective version of the Almgren–Pitts min–max theory due to Nabutovsky–Rotman [20], this yields
$A_{\min}(M,g)\ \leq\ F(v,D)$
for some function $F$ depending only on $v$ and $D$. In that setting, however, the dependence of $f_{1},f_{2}$ and $F$ on the geometric parameters was not made explicit, since the constants coming from the Cheeger–Naber bubble–tree decomposition [8] were only known to exist abstractly.
In the present work we revisit this question in the more rigid Einstein setting. For $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$ the curvature solves an elliptic system, and global Sobolev inequalities can be used, following Anderson [2, 3], to derive quantitative $L^{2}$–curvature bounds and $\varepsilon$–regularity with constants depending only on a Sobolev constant. This analytic input, combined with the Einstein specialization of the Cheeger–Naber bubble–tree decomposition and an improved combinatorial filling estimate, allows us to obtain a quantitative upper bound for $A_{\min}(M,g)$ in terms of $(v,D)$ and the Sobolev constants.
Our first main result is the following quantitative bound for minimal $2$-varifolds in Einstein $4$-manifolds.
Theorem 1.1.
For every $v,D>0$ there exists a constant $F_{\operatorname{Ein}}(v,D)>0$ with the following property. If $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$, then the area $A_{\min}(M,g)$ of a smallest $2$-dimensional stationary integral varifold in $(M,g)$ satisfies
$A_{\min}(M,g)\ \leq\ F_{\operatorname{Ein}}(v,D).$
Moreover, the dependence of $F_{\operatorname{Ein}}(v,D)$ is reduced to the Sobolev constant $C_{S}(v,D)$, the global $L^{2}$-curvature bound, and the standard Einstein $\varepsilon$-regularity constants.
As in [24], the proof of Theorem 1.1 proceeds via a homological filling inequality expressed in terms of the first homological filling function $\operatorname{FH}_{1}$ of $(M,g)$.
Theorem 1.2.
For every $v,D>0$ there exist functions
$f_{1}^{\operatorname{Ein}}(v,D),f_{2}^{\operatorname{Ein}}(v,D)>0,$
with the following property. If $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$ and $C\in\mathcal{Z}_{1}(M;\mathbb{Z})$ is a singular Lipschitz $1$-cycle, then $C$ bounds a singular $2$-chain $E\in\mathcal{C}_{2}(M;\mathbb{Z})$ such that
$\operatorname{mass}_{2}(E)\ \leq\ f_{1}^{\operatorname{Ein}}(v,D)\,\operatorname{mass}_{1}(C)+f_{2}^{\operatorname{Ein}}(v,D).$
Equivalently, the first homological filling function of $(M,g)$ satisfies
$\operatorname{FH}_{1}(l)\ \leq\ f_{1}^{\operatorname{Ein}}(v,D)\,l+f_{2}^{\operatorname{Ein}}(v,D)\qquad\text{for all }l\geq 0.$
Theorem 1.1 is then an immediate consequence of Theorem 1.2 and the min–max theorem of Nabutovsky–Rotman [20], which relates $A_{\min}(M,g)$ to $\operatorname{FH}_{1}$. In the Einstein setting we can track the dependence of the constants arising from the Sobolev inequality, the $L^{2}$-curvature bound and the bubble-tree decomposition, thus reducing the dependence of $F_{\operatorname{Ein}}(v,D)$ to these analytic and combinatorial inputs.
2. Sobolev inequality and Einstein $\varepsilon$-regularity
Throughout this section we assume $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$.
2.1. Global Sobolev inequality
We use the $L^{2}$-Sobolev inequality on $(M,g)$, which is equivalent to a uniform isoperimetric inequality under our assumptions.
Theorem 2.1 (Sobolev inequality).
There exists a constant $C_{S}(v,D)>0$ depending only on $v,D$ such that for every $u\in C^{\infty}(M)$ we have
(2) $\Bigl(\int_{M}|u|^{4}\Bigr)^{1/2}\ \leq\ C_{S}(v,D)\left(\int_{M}|\nabla u|^{2}+\frac{1}{D^{2}}\int_{M}u^{2}\right).$
Proof.
Under our standing assumptions we have $\operatorname{Ric}_{g}\geq-3g$, $\operatorname{Vol}(M)\geq v$ and $\operatorname{diam}(M)\leq D$. A standard consequence is a uniform isoperimetric inequality, hence a uniform Sobolev inequality for mean-zero functions; see, for instance, [16, 2]. Thus there exists $C^{\prime}_{S}(v,D)$ such that
$\|u-\bar{u}\|_{L^{4}(M)}^{2}\leq C^{\prime}_{S}(v,D)\,\|\nabla u\|_{L^{2}(M)}^{2},\qquad\bar{u}:=\operatorname{Vol}(M)^{-1}\int_{M}u.$
In addition, Bishop–Gromov comparison with $\operatorname{Ric}_{g}\geq-3g$ and $\operatorname{diam}(M)\leq D$ gives an upper volume bound $\operatorname{Vol}(M)\leq V_{+}(D)$. Since
$|\bar{u}|^{2}\leq\operatorname{Vol}(M)^{-1}\|u\|_{L^{2}(M)}^{2}\leq v^{-1}\|u\|_{L^{2}(M)}^{2},$
we obtain
$\|\bar{u}\|_{L^{4}(M)}^{2}=|\bar{u}|^{2}\,\operatorname{Vol}(M)^{1/2}\leq v^{-1}V_{+}(D)^{1/2}\,\|u\|_{L^{2}(M)}^{2}\leq C^{\prime\prime}_{S}(v,D)\,\frac{1}{D^{2}}\|u\|_{L^{2}(M)}^{2},$
after enlarging the constant. Combining the bounds for $u-\bar{u}$ and $\bar{u}$ yields (2). ∎
Remark 2.2.
The existence of such a Sobolev constant under a Ricci lower bound, a volume lower bound and a diameter upper bound is classical; see, for example,[16].
$2.$ $L^{2}$ curvature and the curvature equation
We use the local $L^{2}$ curvature estimate of Cheeger–Naber together with a covering argument to obtain a global bound in our setting.
Theorem 2.3 ($L^{2}$ curvature bound).
There exists $C_{L^{2}}(v,D)>0$ such that
$\int_{M}|\operatorname{Rm}_{g}|^{2}\leq C_{L^{2}}(v,D)\qquad\text{for all }(M^{4},g)\in\mathcal{E}_{1}(4,v,D).$
Proof.
Since $|\operatorname{Ric}_{g}|\leq 3$, Bishop–Gromov comparison and the bounds $\operatorname{Vol}(M)\geq v$, $\operatorname{diam}(M)\leq D$ imply that there exists $v_{1}(v,D)>0$ such that
$\operatorname{Vol}(B_{1}(x))\geq v_{1}(v,D)\qquad\text{for every }x\in M.$
Applying the local $L^{2}$ estimate of Cheeger–Naber [8, Theorem 1.13] to each unit ball gives
$\int_{B_{1}(x)}|\operatorname{Rm}|^{2}\leq C_{\mathrm{CN}}(v_{1}(v,D))\qquad\text{for every }x\in M.$
Now choose a maximal $\frac{1}{2}$-separated set $\{x_{1},\dots,x_{N}\}\subset M$. Then the balls $B_{1}(x_{i})$ cover $M$, while the balls $B_{1/4}(x_{i})$ are disjoint. Using again Bishop–Gromov comparison together with the diameter bound, we obtain $N\leq N(v,D)$ for some constant depending only on $(v,D)$. Therefore
$\int_{M}|\operatorname{Rm}|^{2}\leq\sum_{i=1}^{N}\int_{B_{1}(x_{i})}|\operatorname{Rm}|^{2}\leq N(v,D)\,C_{\mathrm{CN}}(v_{1}(v,D)).$
This gives the desired global bound. ∎
Next we recall the elliptic equation satisfied by the curvature tensor. For an Einstein metric $\operatorname{Ric}_{g}=\lambda g$, recall the Weitzenböck formula for $\operatorname{Rm}$:
(3) $\Delta\operatorname{Rm}=\operatorname{Rm}*\operatorname{Rm}+\lambda\,\operatorname{Rm},$
where $\Delta$ is the rough Laplacian and $\operatorname{Rm}*\operatorname{Rm}$ denotes a quadratic expression in $\operatorname{Rm}$; see for example [4]. Let
$u:=|\operatorname{Rm}|.$
Then, using the standard Bochner formula and Kato’s inequality $|\nabla u|\leq|\nabla\operatorname{Rm}|$, one obtains a scalar differential inequality of the form
(4) $\Delta u\ \leq\ C_{0}\,(u^{2}+u)$
for some dimensional constant $C_{0}>0$. We briefly justify this.
Indeed, compute
$\frac{1}{2}\Delta u^{2}=\langle\Delta\operatorname{Rm},\operatorname{Rm}\rangle+|\nabla\operatorname{Rm}|^{2}.$
Using (3) and the Cauchy–Schwarz inequality,
$\langle\Delta\operatorname{Rm},\operatorname{Rm}\rangle=\langle\operatorname{Rm}*\operatorname{Rm},\operatorname{Rm}\rangle+\lambda|\operatorname{Rm}|^{2}\geq-C\,|\operatorname{Rm}|^{3}-C\,|\operatorname{Rm}|^{2}$
for some dimensional constant $C>0$. Therefore
$\frac{1}{2}\Delta u^{2}\geq-Cu^{3}-Cu^{2}+|\nabla\operatorname{Rm}|^{2}.$
On the other hand,
$\frac{1}{2}\Delta u^{2}=u\Delta u+|\nabla u|^{2}.$
Combining these identities and using Kato’s inequality $|\nabla u|\leq|\nabla\operatorname{Rm}|$ yields
$u\Delta u+|\nabla u|^{2}\geq-Cu^{3}-Cu^{2}+|\nabla u|^{2},$
hence
$\Delta u\geq-C(u^{2}+u).$
Rewriting this as $\Delta u\leq C_{0}(u^{2}+u)$ with $C_{0}=C$ gives (4).
2.3. Einstein $\varepsilon$-regularity
The next result is standard in dimension four; we record it in the form needed later.
Theorem 2.4 (Einstein $\varepsilon$-regularity).
There exist constants $\varepsilon_{\mathrm{Ein}}(v,D)>0$, $C_{\mathrm{Ein}}(v,D)>0$ and $c_{\mathrm{Ein}}(v,D)>0$ such that the following holds.
Let $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$ and suppose that $B_{2r}(x)\subset M$ satisfies
$\int_{B_{2r}(x)}|\operatorname{Rm}|^{2}\leq\varepsilon_{\mathrm{Ein}}(v,D).$
Then
$\sup_{B_{r}(x)}|\operatorname{Rm}|\leq C_{\mathrm{Ein}}(v,D)\,r^{-2},\qquad r_{h}(y)\geq c_{\mathrm{Ein}}(v,D)\,r\quad\text{for all }y\in B_{r}(x).$
Proof.
After rescaling we may assume $r=1$. The hypotheses place us in the standard noncollapsed Einstein setting: the Ricci tensor is uniformly bounded, local volume ratios are bounded below in terms of $(v,D)$, and the Sobolev constant is controlled by Theorem 2.1. The usual $\varepsilon$-regularity theorem for Einstein $4$-manifolds therefore applies and yields the pointwise curvature bound together with the harmonic-radius estimate on $B_{1}(x)$; see Anderson [2, 3], Cheeger–Tian [9], and Cheeger–Naber [8]. Rescaling back proves the statement. ∎
Remark 2.5.
In what follows we only use the existence of the constants in Theorems 2.1, 2.3 and 2.4. The global harmonic-radius scale entering the bubble–tree decomposition will be taken directly from Cheeger–Naber’s theorem.
2.4. Bubble–tree decomposition
We state Cheeger–Naber’s structure theorem [8], which will be used to construct a cover on the manifold.
Theorem 2.6 (Bubble–tree decomposition).
There exist constants $r_{0}(v,D)>0$, $C_{\Gamma}(v,D)>0$ and integers $N_{E}(v,D)$, $k_{E}(v,D)$ together with a decomposition
$M=\mathcal{B}^{1}\cup\bigcup_{j_{2}=1}^{N_{2}}\mathcal{N}^{2}_{j_{2}}\cup\bigcup_{j_{2}=1}^{N_{2}}\mathcal{B}^{2}_{j_{2}}\cup\dots\cup\bigcup_{j_{k}=1}^{N_{k}}\mathcal{N}^{k}_{j_{k}}\cup\bigcup_{j_{k}=1}^{N_{k}}\mathcal{B}^{k}_{j_{k}}$
such that:
  1. (i)
    (Bodies) If $x\in\mathcal{B}^{i}_{j}$ then
    $r_{h}(x)\geq r_{0}(v,D)\,\operatorname{diam}(\mathcal{B}^{i}_{j}).$
  2. (ii)
    (Necks) Each $\mathcal{N}^{i}_{j}$ is diffeomorphic to an annulus in $\mathbb{R}^{4}/\Gamma^{i}_{j}$, with $|\Gamma^{i}_{j}|\leq C_{\Gamma}(v,D)$.
  3. (iii)
    Each $\mathcal{N}^{i}_{j}$ meets exactly two bodies, a parent $\mathcal{B}^{i-1}_{\cdot}$ and a child $\mathcal{B}^{i}_{j}$, and their intersections are diffeomorphic to $\mathbb{R}\times S^{3}/\Gamma^{i}_{j}$.
  4. (iv)
    The constants $N_{i}\leq N_{E}(v,D)$ and $k\leq k_{E}(v,D)$.
This is the specialization of Cheeger–Naber’s bubble–tree construction [8, Theorem 8.63] to the Einstein class. The displayed decomposition above is their formula (8.64). We denote the adjacency relations between bodies and necks as a rooted tree with root $\mathcal{B}^{1}$ and denote by $\mathcal{T}^{i}_{j}$ the subtree with root $\mathcal{B}^{i}_{j}$.
3. Filling functions and stationary varifolds
In this section we recall how an upper bound for the first homological filling function controls the area of a smallest $2$-dimensional stationary integral varifold, and we explain how to obtain such an upper bound in the Einstein setting. The overall strategy follows our previous work [24]: we construct a cover adapted to the bubble–tree decomposition, pass to the nerve and an associated geodesic graph, and then fill $1$-cycles by combining local geometric fillings with a combinatorial filling argument.
The one new ingredient in the present paper is an improved combinatorial filling lemma (Lemma 3.7 below), based on a sharp bound for integer solutions of linear Diophantine systems due to Borosh–Flahive–Rubin–Treybig [5] together with Hadamard’s inequality. For the reader’s convenience we include a complete proof of this lemma. All other steps are essentially the same as those in [24], up to the replacement of the general Ricci bounds by the Einstein $\varepsilon$-regularity input from Section 2, and we therefore only sketch the arguments or omit proofs with precise references to [24].
3.1. Homological filling functions and stationary varifolds
Let $\mathcal{C}_{k}(M;\mathbb{Z})$ be the group of singular Lipschitz $k$-chains with integer coefficients and $\mathcal{Z}_{k}(M;\mathbb{Z})=\ker\partial_{k}$ the group of $k$-cycles. If $C=\sum_{i}a_{i}f_{i}\in\mathcal{C}_{k}(M;\mathbb{Z})$ with $f_{i}:\Delta^{k}\to M$ Lipschitz, we define its $k$-mass by
$\operatorname{mass}_{k}(C)\;=\;\sum_{i}|a_{i}|\int_{\Delta^{k}}f_{i}^{*}(d\operatorname{vol}_{g}).$
Assuming $H_{1}(M;\mathbb{Z})=0$, the first homological filling function is defined for $l\geq 0$ by
$\operatorname{FH}_{1}(l)\;=\;\sup_{\operatorname{mass}_{1}(Z)\leq l}\;\inf_{\partial C=Z}\,\operatorname{mass}_{2}(C),$
where the supremum is taken over all singular Lipschitz $1$-cycles $Z\in\mathcal{Z}_{1}(M;\mathbb{Z})$ with $\operatorname{mass}_{1}(Z)\leq l$.
The connection between $\operatorname{FH}_{1}$ and minimal $2$-varifolds is provided by the following theorem of Nabutovsky–Rotman.
Theorem 3.1 (Nabutovsky–Rotman [20]).
Let $M$ be a closed $n$-dimensional Riemannian manifold with $H_{1}(M;\mathbb{Z})=0$ and diameter $\operatorname{diam}(M)=D$. Let $A_{\min}(M)$ denote the area of a smallest $2$-dimensional stationary integral varifold in $M$. Then
$A_{\min}(M)\;\leq\;\frac{(n+1)!}{2}\,\operatorname{FH}_{1}(2D).$
In particular, for $n=4$,
$A_{\min}(M)\;\leq\;60\,\operatorname{FH}_{1}(2D).$
Thus, in order to bound $A_{\min}(M)$ from above, it suffices to obtain a affine upper bound for $\operatorname{FH}_{1}(l)$ in terms of the geometric data of $(M,g)$. The rest of this section is devoted to establishing such a bound in the Einstein setting.
3.2. A homological filling inequality for Einstein manifolds
We now explain the geometric-combinatorial construction that leads to the Einstein filling inequality. The starting point is the bubble–tree decomposition of Theorem 2.6, which decomposes $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$ into finitely many body regions $\mathcal{B}^{i}_{j}$ and neck regions $\mathcal{N}^{i}_{j}$, with quantitative control on harmonic radius in the bodies and on the topology and geometry of the necks.
The idea is to build:
  • (a)
    a cover of $M$ by harmonic balls inside bodies and by “trapezoids” inside necks, with uniformly controlled overlap,
  • (b)
    the nerve $\mathfrak{N}$ of this cover and a canonical map $f:M\to\mathfrak{N}$, and
  • (c)
    a geodesic graph $\Gamma\subset M$ whose vertices correspond to the elements of the cover and whose edges have uniformly bounded length.
A general $1$-cycle $C$ in $M$ will first be homologically replaced by a sum of cycles supported in single bodies, then approximated by a cycle in $\Gamma$, then straightened via the standard graph-to-nerve procedure of [24, Sec. 4] to a simplicial cycle in $\mathfrak{N}$ and filled combinatorially using Lemma 3.7. Each $2$-simplex in $\mathfrak{N}$ corresponds to a geodesic triangle in $\Gamma$, which can be filled by a $2$-chain of uniformly bounded area. This yields the desired affine filling inequality.
Bodies. For each body $\mathcal{B}^{i}_{j}$, let
$r^{i}_{j}:=\inf_{x\in\mathcal{B}^{i}_{j}}r_{h}(x)\;\geq\;r_{0}(v,D)\,\operatorname{diam}(\mathcal{B}^{i}_{j}),$
where $r_{h}$ is the harmonic radius and $r_{0}(v,D)$ is the constant from Theorem 2.6. Set
$R^{i}_{j}:=\frac{r^{i}_{j}}{64}.$
By volume comparison and the lower bound on $r^{i}_{j}$, one can cover $\mathcal{B}^{i}_{j}$ by at most $N_{1}(v,D)$ balls of radius $R^{i}_{j}$, and we denote a fixed such subcover by $\mathcal{O}(\mathcal{B}^{i}_{j})$.
Necks. For each neck $\mathcal{N}^{i}_{j}\simeq A_{\rho,2\rho}(0)\subset\mathbb{R}^{4}/\Gamma^{i}_{j}$ via a diffeomorphism $\Phi^{i}_{j}$, we cover the spherical factor $S^{3}/\Gamma^{i}_{j}$ by $m_{i,j}\leq N_{2}(v,D)$ convex balls $B_{r_{c}/4}(z_{\ell})$ and define “trapezoids”
$T_{i,j,\ell}=\Phi^{i}_{j}\bigl((2\sqrt{1-\varepsilon}\,\rho,(2-\sqrt{1-\varepsilon})\,\rho)\times B_{r_{c}/4}(z_{\ell})\bigr),$
contained in slightly larger convex regions
$\bar{T}_{i,j,\ell}=\Phi^{i}_{j}\bigl((\rho/2,2\rho)\times B_{r_{c}}(z_{\ell})\bigr),$
where $\varepsilon>0$ and $r_{c}>0$ are chosen (as in [24]) so that the $\bar{T}_{i,j,\ell}$ are contained in the neck and have controlled overlap.
We then define the global cover
$\mathcal{O}:=\Big(\bigcup_{i,j}\mathcal{O}(\mathcal{B}^{i}_{j})\Big)\;\cup\;\Big(\bigcup_{i,j,\ell}T_{i,j,\ell}\Big).$
The number of sets in $\mathcal{O}$ is bounded by $|\mathcal{O}|\leq\tilde{N}_{E}(v,D)$ for some $\tilde{N}_{E}(v,D)$ depending only on the bubble–tree complexity $N_{E}(v,D)$, $k_{E}(v,D)$ and the covering constants $N_{1}(v,D)$, $N_{2}(v,D)$; see also [24, Sec. 4] for the general Ricci-bounded case.
Let $\mathfrak{N}$ be the nerve of $\mathcal{O}$, and let $f:M\to\mathfrak{N}$ denote the canonical map associated to a fixed partition of unity subordinate to $\mathcal{O}$.
Geodesic graph $\Gamma$. We now construct a geodesic graph $\Gamma\subset M$ whose vertices correspond to the centers of the body balls in $\mathcal{O}(\mathcal{B}^{i}_{j})$ and to chosen reference points in the trapezoids $T_{i,j,\ell}$. Whenever two elements of $\mathcal{O}$ intersect, we connect the corresponding vertices in one of the following ways:
  • (i)
    if both sets lie in the same body, we join the centers by a minimizing geodesic inside that body;
  • (ii)
    if both sets are trapezoids in the same neck, we join the reference points by a short curve contained in the union of the two trapezoids;
  • (iii)
    if one set is a body ball and the other is a trapezoid in a neighbouring neck, we concatenate a short geodesic segment inside the ball with a short curve inside the trapezoid.
By construction, each edge of $\Gamma$ has length bounded by
$\leq C(v,D)\,\max\{R^{i}_{j},\rho\},$
for some constant $C(v,D)$ depending only on the geometric data. This construction is identical to that in [24, Sec. 4], with the only difference that the scales $R^{i}_{j}$ and $\rho$ are now tied to the harmonic-radius scale $r_{0}(v,D)$ furnished by Theorem 2.6.
3.3. Local fillings
We next recall the local geometric fillings that will be used to fill the small pieces of a $1$-cycle inside bodies and necks.
Lemma 3.2.
In a harmonic chart $\Phi:B_{r_{h}}(0)\to M$ with
$\|g-\delta\|_{C^{0}}+r_{h}\|\partial g\|_{C^{0}}\leq 10^{-3},$
any curve $\gamma$ satisfies
$\sqrt{1-10^{-3}}\;\operatorname{length}(\gamma)\;\leq\;\operatorname{length}(\Phi\gamma)\;\leq\;\sqrt{1+2\cdot 10^{-3}}\;\operatorname{length}(\gamma),$
and $\Phi(B_{r_{h}}(0))\supset B_{r_{h}/2}(\Phi(0))$.
Lemma 3.3.
Let $r_{h}(x)>0$ and let $C$ be a $1$-cycle whose image is contained in $B_{r_{h}(x)/2}(x)$. Then there exists a $2$-chain $E$ with $\partial E=C$ and
$\operatorname{mass}_{2}(E)\;\leq\;r_{h}(x)\,\operatorname{mass}_{1}(C).$
Proof.
In harmonic coordinates $\Phi:B_{r_{h}}\to M$, the metric is $(1\pm 10^{-3})$-bilipschitz to the Euclidean metric by Lemma 3.2. Coning $\Phi^{-1}(C)$ to 0 in $\mathbb{R}^{4}$ produces a $2$-chain of Euclidean area $\leq\frac{1}{2}r_{h}\,\operatorname{length}(\Phi^{-1}C)$. Pushing forward via $\Phi$ distorts area by at most a fixed factor close to $1$, which we absorb into the constant. ∎
Lemma 3.4 (Neck contraction).
Let $\mathcal{N}\simeq A_{\rho,2\rho}(0)\subset\mathbb{R}^{4}/\Gamma$ be a neck region in the bubble–tree decomposition, and let $C$ be a $1$-cycle with image in $\mathcal{N}$. Then there exist a $1$-cycle $C^{\prime}$ supported in $\mathcal{N}\cap\mathcal{B}$ (the adjacent child body) and a $2$-chain $E$ such that
$\partial E=C-C^{\prime},\ \operatorname{mass}_{1}(C^{\prime})\leq h_{E}(v,D)\,\operatorname{diam}(\mathcal{B}),\ \operatorname{mass}_{2}(E)\leq C(v,D)\,\rho\,\operatorname{mass}_{1}(C),$
where $h_{E}(v,D)$ depends only on the uniform bound $C_{\Gamma}(v,D)$ from Theorem 2.6.
Proof.
The argument is exactly as in [24, Sec. 4.2]. Via $\Phi^{-1}$ we view $C$ as a $1$-cycle in the annulus $A_{\rho,2\rho}(0)\subset\mathbb{R}^{4}/\Gamma$. Radial projection onto a sphere $S_{\lambda}/\Gamma$ with $\lambda\in(\rho,2\rho)$ sweeps out a $2$-chain of area $\lesssim\rho\,\operatorname{mass}_{1}(C)$. The projected cycle is a $1$-cycle in $S^{3}/\Gamma$, whose first homology is finite with generators represented by loops of length $\leq 2\pi\lambda$. Replacing the projected cycle by a representative of its homology class supported in a bounded number of such generators yields $C^{\prime}$ with $\operatorname{mass}_{1}(C^{\prime})\leq h_{E}(v,D)\,\lambda$. Pushing forward by $\Phi$ and controlling the distortion inside the annulus completes the proof. ∎
3.4. Approximation by $\Gamma$ and improved combinatorial filling
The next step is to replace a $1$-cycle lying in a body region by a combinatorial cycle in the geodesic graph $\Gamma$.
Lemma 3.5.
Let $C$ be a $1$-cycle supported in a body region $\mathcal{B}$. Then there exists a $1$-cycle $C^{\prime}\subset\Gamma$ and a $2$-chain $E$ such that
$\partial E=C-C^{\prime},\ \operatorname{mass}_{1}(C^{\prime})\leq C_{\Gamma}(v,D)\,\operatorname{mass}_{1}(C),\ \operatorname{mass}_{2}(E)\leq C_{\Gamma}(v,D)\,\operatorname{mass}_{1}(C).$
Proof.
The proof is identical to [24, Lemma 4.5]. One partitions $C$ into short arcs contained in the balls of the cover $\mathcal{O}(\mathcal{B})$, with length $\leq R^{i}_{j}$. Inside each ball one joins the endpoints of an arc to the corresponding center along edges of $\Gamma$. The difference between the original arc and the resulting path in $\Gamma$ lies in a harmonic ball and hence can be filled using Lemma 3.3. Summing over all arcs gives the desired estimates. ∎
The idea of controlling the combinatorial filling via integer solutions of linear systems was introduced in [20]. We show that one can obtain an upper bound for the number of simplices needed to fill a simplicial cycle which depends linearly on the size of the cycle. Note that this is independent of the dimension of the simplicial complex.
Let $A\in\mathbb{Z}^{\ell\times n}$, $b\in\mathbb{Z}^{\ell}$, and assume $\mathrm{rank}(A)=m\leq\ell$. Set
$M_{A}:=\max_{1\leq i\leq\ell,\;1\leq j\leq n}|a_{ij}|,\qquad M_{b}:=\max_{1\leq i\leq\ell}|b_{i}|.$
Let $Y$ be the maximum of the absolute values of all $m\times m$ minors of the augmented matrix (A|b)(A\,|\,b):
$Y:=\max\bigl\{\,|\det S|:\ S\text{ is an $m\times m$ submatrix of }(A\,|\,b)\,\bigr\}.$
Theorem 3.6.
Assume that the system $Ax=b$ has an integer solution. Then there exists an integer solution $x=(x_{i})\in\mathbb{Z}^{n}$ such that
(5) $\max_{i}|x_{i}|\;\leq\;m^{m/2}\,\max\{\,M_{A}^{m},\;M_{A}^{m-1}M_{b}\,\}\;=\;m^{m/2}M_{A}^{m-1}\max\{M_{A},M_{b}\}.$
In particular, this upper bound depends only on $m$, $M_{A}$ and $M_{b}$ (and is independent of $n$).
Proof.
By [5, Thm. 1], if $Ax=b$ has an integer solution then there exists an integer solution $x$ with
(6) $\max_{i}|x_{i}|\;\leq\;Y,$
where $Y$ is the maximum of the absolute values of the $m\times m$ minors of (A|b)(A\,|\,b) defined above.
We estimate $Y$ using Hadamard’s inequality. Let $S$ be any $m\times m$ submatrix of (A|b)(A\,|\,b) and denote its columns by $s_{1},\dots,s_{m}\in\mathbb{R}^{m}$. Hadamard’s inequality gives
(7) $|\det S|\;\leq\;\prod_{k=1}^{m}\|s_{k}\|_{2}.$
There are two cases.
Case 1: $S$ is formed entirely by columns of $A$. Then every entry of $S$ has absolute value at most $M_{A}$, hence each column has Euclidean norm $\|s_{k}\|_{2}\leq\sqrt{m}\,M_{A}.$ By (7),
$|\det S|\;\leq\;(\sqrt{m}M_{A})^{m}\;=\;m^{m/2}M_{A}^{m}.$
Case 2: $S$ contains the column $b$. Then one column of $S$ (the one coming from $b$) has norm at most $\sqrt{m}\,M_{b}$, while each of the remaining m1m-1 columns has norm at most $\sqrt{m}\,M_{A}$. Thus
$|\det S|\;\leq\;(\sqrt{m}M_{b})\,(\sqrt{m}M_{A})^{m-1}\;=\;m^{m/2}M_{A}^{m-1}M_{b}.$
Taking the maximum over all $S$ in these two cases yields
$Y\;\leq\;m^{m/2}\,\max\{M_{A}^{m},\;M_{A}^{m-1}M_{b}\}.$
Combining this with (6) gives (5). ∎
Lemma 3.7.
Let $X$ be an $n$-dimensional simplicial complex with $n_{0}$ vertices. Let
$c_{k}=\sum_{i}a_{k}^{i}\,\sigma_{i}^{k}$
be a simplicial $k$-boundary in $X$ with integer coefficients. Then there exists a simplicial $(k+1)$-chain
$c_{k+1}=\sum_{i}a_{k+1}^{i}\,\sigma_{i}^{k+1}$
such that $\partial c_{k+1}=c_{k}$ and
(8) $\max_{i}|a_{k+1}^{i}|\;\leq\;\binom{n_{0}}{k+1}^{\frac{1}{2}\binom{n_{0}}{k+1}}\,\max_{i}|a_{k}^{i}|.$
In particular, since the number $n_{k+1}$ of $(k+1)$-simplices satisfies $n_{k+1}\leq\binom{n_{0}}{k+2}$, we also have
(9) $\sum_{i}|a_{k+1}^{i}|\;\leq\;\binom{n_{0}}{k+2}\,\binom{n_{0}}{k+1}^{\frac{1}{2}\binom{n_{0}}{k+1}}\,\max_{i}|a_{k}^{i}|.$
Proof.
If $c_{k}=0$, the conclusion is trivial with $c_{k+1}=0$, so we assume $c_{k}\neq 0$. Let $C_{j}$ denote the group of $j$-chains of $X$ with integer coefficients. We fix the usual basis of $C_{j}$ given by all oriented $j$-simplices $\sigma^{j}_{1},\dots,\sigma^{j}_{n_{j}}$, where $n_{j}$ is the number of $j$-simplices of $X$.
Consider the boundary map
$\partial_{k+1}\colon C_{k+1}\to C_{k}$
and write its matrix with respect to these bases as $A\in\mathbb{Z}^{n_{k}\times n_{k+1}}.$ By construction, each entry of $A$ lies in $\{0,\pm 1\}$, so
$M_{A}:=\max_{i,j}|a_{ij}|=1.$
Write $c_{k}$ in coordinates as $c_{k}=\sum_{i}a_{k}^{i}\sigma_{i}^{k},$ and let $b\in\mathbb{Z}^{n_{k}}$ be the column vector with entries $b_{i}=a_{k}^{i}$. Then the condition $\partial c_{k+1}=c_{k}$ is exactly the linear system
$Ax=b,$
where $x\in\mathbb{Z}^{n_{k+1}}$ is the coordinate vector of $c_{k+1}$. Since $c_{k}$ is a boundary, this system has at least one integer solution. Set
$M_{b}:=\max_{i}|b_{i}|=\max_{i}|a_{k}^{i}|\ (\geq 1).$
Let $m:=\mathrm{rank}(A)\leq n_{k}$. Applying Theorem 3.6 with this $A$ and $b$, and using $M_{A}=1$, we obtain an integer solution $x$ of $Ax=b$ such that
(10) $\max_{i}|x_{i}|\;\leq\;m^{m/2}\,\max\{M_{A},\;M_{b}\}\;=\;m^{m/2}\,\max_{i}|a_{k}^{i}|.$
Define $c_{k+1}:=\sum_{i}x_{i}\sigma_{i}^{k+1}$; then $\partial c_{k+1}=c_{k}$, and the coefficients $a_{k+1}^{i}$ of $c_{k+1}$ are exactly the $x_{i}$.
Next we bound $m$ in terms of the number of vertices $n_{0}$. Any $k$-simplex of $X$ is determined by its $(k+1)$ vertices, so the number of $k$-simplices satisfies
$n_{k}\leq\binom{n_{0}}{k+1}$.
Hence
$m=\mathrm{rank}(A)\leq n_{k}\leq\binom{n_{0}}{k+1}$.
Combining this with (10) yields
$\max_{i}|a_{k+1}^{i}|=\max_{i}|x_{i}|\;\leq\;m^{m/2}\,\max_{i}|a_{k}^{i}|\;\leq\;\binom{n_{0}}{k+1}^{\frac{1}{2}\binom{n_{0}}{k+1}}\max_{i}|a_{k}^{i}|,$
which is exactly (8). Finally, since there are at most $n_{k+1}\leq\binom{n_{0}}{k+2}$ nonzero coefficients $a_{k+1}^{i}$, we also have
$\sum_{i}|a_{k+1}^{i}|\;\leq\;n_{k+1}\,\max_{i}|a_{k+1}^{i}|\;\leq\;\binom{n_{0}}{k+2}\,\binom{n_{0}}{k+1}^{\frac{1}{2}\binom{n_{0}}{k+1}}\,\max_{i}|a_{k}^{i}|,$
which is (9). ∎
In our application we will only use Lemma 3.7 for $k=1$ and for the simplicial complex $X=\mathfrak{N}$ (the nerve of $\mathcal{O}$).
Lemma 3.8 (Filling geodesic triangles).
There exists a constant $K_{9}(v,D)>0$ such that any geodesic triangle formed by three edges of $\Gamma$ bounds a $2$-chain $E$ with
$\operatorname{mass}_{2}(E)\leq K_{9}(v,D).$
Proof.
As in [24, Lemma 4.6]. If all three vertices lie inside a single body, the triangle lies in a harmonic ball $B_{r_{h}/2}$ and can be filled using Lemma 3.3. If they lie inside a single neck, the triangle lies in the union of at most two trapezoids contained in $\bar{T}_{i,j,\ell}$, and can be filled by a cone in the annulus, with area $\lesssim\rho^{2}$. Mixed configurations are reduced to these two cases. All constants can be controlled in terms of $r_{0}(v,D)$ and $D$. ∎
3.5. Decomposition and the Einstein filling inequality
We now put the pieces together. The first step is to decompose a general $1$-cycle according to the bubble–tree structure.
Lemma 3.9 (Decomposition along the tree).
Let $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$ and let $C\in\mathcal{Z}_{1}(M;\mathbb{Z})$ be a singular Lipschitz $1$-cycle. Then there exists a decomposition
$C=\sum_{i=1}^{N}C_{i},$
with $N\leq N_{E}(v,D)\,k_{E}(v,D)$, such that each $C_{i}$ is supported in a single body or a single neck and
$\sum_{i=1}^{N}\operatorname{mass}_{1}(C_{i})\;\leq\;C_{0}(v,D)\,\operatorname{mass}_{1}(C).$
Proof.
As in [24, Sec. 4.3], one cuts $C$ whenever it crosses a boundary component $\partial\mathcal{B}\cap\mathcal{N}$ and reassembles the resulting arcs into cycles supported in individual bodies or necks. The number of such regions is bounded by the bubble–tree complexity $N_{E}(v,D)\,k_{E}(v,D)$, and the total length increases by at most a uniform multiplicative factor $C_{0}(v,D)$, coming from bounded overlap of the decomposition and the uniform control on the diameter of bodies and necks. ∎
Corollary 3.10 (Reduction to bodies).
Let $C$ be as in Lemma 3.9. Then $C$ is homologous to a $1$-cycle $C^{\mathrm{body}}$ supported in the union of body regions, i.e. there exists a $2$-chain $E_{0}$ such that
$\partial E_{0}=C-C^{\mathrm{body}},$
and
$\operatorname{mass}_{1}(C^{\mathrm{body}})\leq C_{1}(v,D)\,\operatorname{mass}_{1}(C),\ \operatorname{mass}_{2}(E_{0})\leq C_{2}(v,D)\,\operatorname{mass}_{1}(C)+C_{3}(v,D).$
Proof.
Apply Lemma 3.4 to each neck-supported component $C_{i}$ in the decomposition of Lemma 3.9. This replaces each $C_{i}$ by a cycle supported in the adjacent body at the expense of adding a $2$-chain whose mass is controlled by $C(v,D)\,\rho\,\operatorname{mass}_{1}(C_{i})$. Summing over all necks and using the uniform bounds on the number and geometry of necks yields the stated estimates. ∎
We are now ready to state and prove the main filling inequality in the Einstein class.
Theorem 3.11 (Einstein filling inequality).
There exist $f_{1}^{\mathrm{Ein}}(v,D),f_{2}^{\mathrm{Ein}}(v,D)>0$ such that for any $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$ and any $C\in\mathcal{Z}_{1}(M;\mathbb{Z})$ there exists $E\in\mathcal{C}_{2}(M;\mathbb{Z})$ with
$\partial E=C,\ \operatorname{mass}_{2}(E)\leq f_{1}^{\mathrm{Ein}}(v,D)\,\operatorname{mass}_{1}(C)+f_{2}^{\mathrm{Ein}}(v,D).$
Equivalently, the first homological filling function satisfies
$\operatorname{FH}_{1}(l)\leq f_{1}^{\mathrm{Ein}}(v,D)\,l+f_{2}^{\mathrm{Ein}}(v,D)\ \text{for all }l\geq 0.$
Proof.
By Corollary 3.10, $C$ is homologous to a cycle $C^{\mathrm{body}}$ supported in the union of bodies, with
$\partial E_{0}=C-C^{\mathrm{body}},\quad\operatorname{mass}_{2}(E_{0})\leq C_{2}(v,D)\,\operatorname{mass}_{1}(C)+C_{3}(v,D),$
and $\operatorname{mass}_{1}(C^{\mathrm{body}})\leq C_{1}(v,D)\,\operatorname{mass}_{1}(C).$
Next, approximate $C^{\mathrm{body}}$ by a cycle in the geodesic graph $\Gamma$ using Lemma 3.5. We obtain a cycle $\widehat{C}\subset\Gamma$ and a $2$-chain $E_{1}$ such that
$\partial E_{1}=C^{\mathrm{body}}-\widehat{C},\ \operatorname{mass}_{2}(E_{1})\leq C_{\Gamma}(v,D)\,\operatorname{mass}_{1}(C^{\mathrm{body}})\leq C_{\Gamma}(v,D)\,C_{1}(v,D)\,\operatorname{mass}_{1}(C).$
The remaining step is the chain-level passage from the graph $\Gamma$ to the nerve $\mathfrak{N}$. Instead of using the partition-of-unity map directly on $\widehat{C}$, we invoke the graph-to-nerve straightening procedure from [24, Sec. 4]. Applied to the graph cycle $\widehat{C}$, it yields a simplicial $1$-boundary
$c_{1}=\sum_{i}a_{1}^{i}\,\sigma_{i}^{1}$
in $\mathfrak{N}$ and a $2$-chain $E_{\mathrm{st}}$ in $M$ such that
$\partial E_{\mathrm{st}}=\widehat{C}-\gamma(c_{1}),$
where $\gamma(c_{1})$ denotes the edge cycle in $\Gamma$ corresponding to $c_{1}$, and
$\sum_{i}|a_{1}^{i}|\leq C_{L}(v,D)\,\operatorname{mass}_{1}(C),\qquad\operatorname{mass}_{2}(E_{\mathrm{st}})\leq C_{L}(v,D)\,\operatorname{mass}_{1}(C).$
In particular,
$\max_{i}|a_{1}^{i}|\leq C_{L}(v,D)\,\operatorname{mass}_{1}(C).$
Let $n_{0}$ be the number of vertices of $\mathfrak{N}$, so $n_{0}\leq\tilde{N}_{E}(v,D)$. Since $c_{1}$ is a simplicial boundary in $\mathfrak{N}$, Lemma 3.7 with $k=1$ yields a simplicial $2$-chain $C_{2}=\sum a_{2}^{i}\sigma_{i}^{2}$ with $\partial C_{2}=c_{1}$ and
$\sum_{i}|a_{2}^{i}|\;\leq\;B_{1}(n_{0})\,\max_{i}|a_{1}^{i}|\;\leq\;B_{1}(\tilde{N}_{E}(v,D))\,C_{L}(v,D)\,\operatorname{mass}_{1}(C),$
where $B_{1}(n_{0})$ is the combinatorial coefficient from (9). Each $2$-simplex in $\mathfrak{N}$ corresponds to a geodesic triangle in $\Gamma$, which bounds a $2$-chain in $M$ of area at most $K_{9}(v,D)$ by Lemma 3.8. Thus $C_{2}$ lifts to a $2$-chain $E_{2}$ in $M$ with
$\partial E_{2}=\gamma(c_{1}),\ \operatorname{mass}_{2}(E_{2})\leq K_{9}(v,D)\,\sum_{i}|a_{2}^{i}|\leq K_{9}(v,D)\,B_{1}(\tilde{N}_{E}(v,D))\,C_{L}(v,D)\,\operatorname{mass}_{1}(C).$
Finally, set
$E:=E_{0}+E_{1}+E_{\mathrm{st}}+E_{2}.$
Then $\partial E=C$, and combining the three estimates we obtain
$\operatorname{mass}_{2}(E)\leq\bigl(C_{2}+C_{\Gamma}C_{1}+C_{L}+K_{9}B_{1}(\tilde{N}_{E})C_{L}\bigr)\operatorname{mass}_{1}(C)\;+\;C_{3}.$
We may therefore take
$f_{1}^{\mathrm{Ein}}(v,D):=C_{2}+C_{\Gamma}C_{1}+C_{L}+K_{9}\,B_{1}(\tilde{N}_{E})\,C_{L},$
and
$f_{2}^{\mathrm{Ein}}(v,D):=C_{3}(v,D).$
[…]
Acknowledgements
The author Z. Zhu was supported in part by National Key R&D Program of China 2024YFA1015300 and NSFC 1250010531.